A manufacturer knows that the razor blades he makes contain on an average 0.5% of defectives. he packs them in packs of 5 what's the probability that a packet picked at random will contain 3 or more faulty blades ?
1
Expert's answer
2020-01-27T09:11:04-0500

Probability of defect p = 0.5% = 0.005

q = 1 - p = 1 - 0.005 = 0.995

Now the number of packs is n = 5.

We need to find probability of 3 or more faulty blades.


"P(x \geq 3)=P(x=3)+P(x=4)+P(x=5)" P(x >= 3)

= 5C3 0.0053 * 0.9952 + 5C4 0.0054*0.995 + 5C5 * 0.0055

= 0.000001238 + 0.000000003 + 3.125 * 10-12

= 1.241 * 10-6